OpenStudy (anonymous):

If f(x) = x2 − 2x + 9 and g(x) = 8 − x, what is (f o g)(−4)?

OpenStudy (anonymous):

121

OpenStudy (whpalmer4):

@loop_51 might want to check your work...

OpenStudy (whpalmer4):

\[f(x) = x^2-2x+9\]\[g(x) = 8-x\]To make \((f\circ g)(x)\), we write out \(f(x)\) and replace \(x\) with the value of \(g(x)\) everywhere we see it. \[(f\circ g)(x) = (8-x)^2-2(8-x)+9\]Now plug in \(x=-4\) and evaluate.

OpenStudy (anonymous):

129?

OpenStudy (whpalmer4):

That's what I get.

OpenStudy (anonymous):

yah my bad 129